A field of E=100 v/m is kept in copper. Determine the power loss per volume



1
Expert's answer
2021-05-16T18:00:44-0400

"\omega=\frac{\delta Q}{dV\cdot dt}=E^2/\rho=100^2/1.75\cdot10^{-8}=5.7\cdot10^{12} \ (J/(m^3\cdot s))"